# Train for the Worst, Plan for the Best: Understanding Token Ordering in Masked Diffusions

### v1

@@ 1 Introduction
While diffusion models (Ho et al., 2020; Song et al., 2021) are now the dominant approach for generative modeling in continuous domains like image, video, and audio, efforts to extend this methodology to discrete domains like text and proteins (Austin et al., 2021; Lou et al., 2024; Hoogeboom et al., 2021) remain nascent. Among numerous proposals, masked diffusion models (MDMs) (Lou et al., 2024; Sahoo et al., 2024; Shi et al., 2024) have emerged as a leading variant, distinguished by a simple and principled objective: to generate samples, learn to reverse a noise process which independently and randomly masks tokens.
In many applications, such as language modeling, masked diffusion models (MDMs) still underperform compared to autoregressive models (ARMs) (Nie et al., 2024; Zheng et al., 2024a), which instead learn to reverse a noise process that unmasks tokens sequentially from left to right. However, recent studies suggest that MDMs may offer advantages in areas where ARMs fall short, including reasoning (Nie et al., 2024; Kitouni et al., 2024), planning (Ye et al., 2024), and infilling (Gong et al., 2024). This raises a key question: what are the strengths and limitations of MDMs compared to ARMs, and under what conditions can MDMs be scaled to challenge the dominance of ARMs in discrete generative modeling?
@@
  • Flexibility at inference time: On the other hand, the sampling paths taken by an MDM are less rigid. The order in which tokens are decoded at inference time is random instead of fixed to left-to-right. In fact, even more is possible: MDMs can actually be used to decode in any order (Zheng et al., 2024a).
@@
While the above might appear to be bad news for MDMs, in the second part of this paper we answer our guiding question in the affirmative by building upon the observation (Zheng et al., 2024a) that MDMs which can perfectly solve all masking subproblems can be used to decode in any order.
@@ 2 Masked Diffusion Models (MDM)
In this section, we explain the framework of Masked Diffusion Models (Shi et al., 2024; Sahoo et al., 2024) and its interpretation as an order-agnostic learner. MDMs gradually add noise to the true discrete data and learn the marginal distribution of the induced reverse process. Below, we formulate the forward and reverse processes for MDMs.
@@
The reverse transition probability $q_{s|t}(x_{s}^{i}|x_{t},x_{0})$ is approximated using $g_{\theta}(x_{s}^{i}|x_{t})\triangleq q_{s|t}(x_{s}^{i}\,\lvert\,x_{t},x_{0}\leftarrow p_{\theta}(x_{t},t))$ where $p_{\theta}(x_{t},t)$ is a denoising network trained to predict the marginal on $x_{0}$ via an ELBO-based loss:
@@ 2.1.1 Order-agnostic training of MDMs
Recent works (Zheng et al., 2024a; Ou et al., 2024) have observed that the learning problem of MDM is equivalent to a masked language model. Building upon their analysis, we reformulate the loss $\mathcal{L}_{\theta}$ to show that $\mathcal{L}_{\theta}$ is a linear combination of the loss for all possible infilling masks. We first define $x_{0}[M]$ as a masked sequence, obtained from original sequence $x_{0}$ where indices in the mask set $M$ (regarded as a subset of $[L]\triangleq\{1,2,\ldots,L\}$) are replaced with mask token 00.
@@ 3.2 Empirical evidence of hardness via likelihoods
Recent studies (Nie et al., 2024; Zheng et al., 2024a) have shown that masked diffusion models (MDMs) underperform compared to autoregressive models (ARMs) on natural text data. In this section, we provide evidence that this performance gap is primarily due to the order-agnostic training of MDMs. Since natural text follows a left-to-right token order, we demonstrate that as training deviates from this order, model performance gradually deteriorates.
@@ 4.1 Effective design of ordering oracle
Top-$K$ probability (Zheng et al., 2024b). Suppose we want to unmask $K$ positions at time step $t$, i.e., select $|S|=K$. In the Top-$K$ strategy, the uncertainty of a position is estimated by the maximum probability assigned to any value in the vocabulary. More precisely, the certainty at position $i$ is $\max_{j\in\{0,\ldots,m-1\}}p_{\theta}(x^{i}=j|x_{t})$ and $\mathcal{F}(\theta,x_{t})=\text{Top }K\left(\max p_{\theta}(x^{i}|x_{t})\right)$.
Top-$K$ strategy is a good proxy for many tasks and works well in practice (Zheng et al., 2024b; Ye et al., 2024; Wang et al., 2024). However, this approach can often provide misleading estimates of uncertainty. Consider when an MDM is confused between two token values, thus assigning them almost equal but high probabilities. In this case, Top-$K$ strategy may still choose to unmask this position, despite its uncertainty. To mitigate this issue, we propose the following alternative strategy.
@@ 4.4 Easy to hard generalization
We keep the training dataset the same as proposed in Shah et al. (2024). Shah et al. (2024) created this dataset from Radcliffe (2020) by filtering out the puzzles that can be solved using 7 fixed strategies and do not require backtracking-based search. We use the remaining puzzles in Radcliffe (2020) as our hard dataset. Hence, these puzzles all use a strategy not seen during training and/or backtracking to obtain the correct solution.
@@
Method ##\##Parameters Accuracy
@@ Appendix A Related works
(Continuous) diffusion models were originally built on continuous-space Markov chains with Gaussian transition kernels (Sohl-Dickstein et al., 2015; Ho et al., 2020). This was later extended to continuous time through the theory of stochastic differential equations (Song et al., 2021). In a similar vein, discrete diffusion models have emerged from discrete-space Markov chains (Hoogeboom et al., 2021). Specifically, (Austin et al., 2021) introduced D3PM with various types of transition matrices. Later, Lou et al. (2024) proposed SEDD, incorporating a theoretically and practically robust score-entropy objective. Additionally, Varma et al. (2024); Liu et al. (2024b) introduced novel modeling strategies that classify tokens in a noisy sequence as either signal (coming from clean data) or noise (arising from the forward process). In particular, Liu et al. (2024b) uses this to give a planner that adaptively determines which tokens to denoise. While this is similar in spirit to our general discussion about devising adaptive inference strategies, we emphasize that their approach is specific to discrete diffusions for which the forward process scrambles the token values, rather than masking them.
@@
Meanwhile, the absorbing transition kernel has gained popularity as a common choice due to its better performance than other kernels. Building on this, Sahoo et al. (2024); Shi et al. (2024) aligned its framework with continuous diffusion, resulting in a simple and principled training recipe, referring to it as Masked Diffusion Model. Subsequent studies have explored various aspects of MDM. Gong et al. (2024) efficiently trained MDM via adaptation from autoregressive models, scaling MDM up to 7B parameters. Zheng et al. (2024a) interpreted MDMs as order-agnostic learners and proposed a first-hitting sampler based on this insight. Ye et al. (2024); Gong et al. (2024) demonstrated that MDM outperforms autoregressive models in reasoning and planning tasks, emphasizing its impact on downstream applications. Nie et al. (2024) examined the scaling laws of MDM, while Xu et al. (2024); Liu et al. (2024a) identified limitations in capturing coordinate dependencies when the number of sampling steps is small and proposed additional modeling strategies to address this issue. Schiff et al. (2024) studied conditional generation using MDM and Rector-Brooks et al. (2024) tackled the challenge of controlling generated data distributions through steering methodologies. Chen & Ying (2024) provided a theoretical analysis showing that sampling error is small given accurate score function estimation.
@@
Even though language tasks generally have a natural order of “left-to-right” token generation, in many tasks like planning, reasoning, and combinatorial optimization, the natural order of token generation can be quite different from “left-to-right”. Even though prominent autoregressive-based language models achieve impressive performance on various tasks, many works (Golovneva et al., 2024; Chen et al., 2024; Kitouni et al., 2024) have shown that this performance is tied to the training order of the tasks and therefore can cause brittleness from it. For example, Chen et al. (2024) showed that simply permuting the premise order on math tasks causes a performance drop of 30%. The reason behind such brittleness regarding the ordering is the inherent “left-to-right” nature of the autoregressive models. Several works (Liao et al., 2020) have tried to address this issue in the autoregressive framework. In particular, (Papadopoulos et al., 2024) highlighted the significance of left-to-right ordering in natural language by comparing its likelihood to that of the reverse (right-to-left) ordering.
Recently, discrete diffusion models have emerged as a promising approach for discrete data apart from autoregressive models. Additionally, the order-agnostic training of discrete diffusion models opens up the multiple sampling paths during the inference but it also faces some challenges during the training therefore, they seem a promising approach to elicit any order reasoning. Zheng et al. (2024b) proposed different ways of implementing an adaptive inference strategy for MDM but a concrete understanding of why such an adaptive inference strategy is needed is still lacking. In this work, we explore various aspects of vanilla MDM training and how adaptive MDM inference can mitigate the issues raised by vanilla MDM training and elicit any order reasoning.
@@
Efforts to learn the natural language using non-autoregressive modeling began with BERT (Devlin et al., 2019). Non-causal approaches can take advantage of the understanding the text data representation. (Chang et al., 2022) adopted a similar approach for learning image representations. Building on these intuitions, (Shih et al., 2022; Hoogeboom et al., 2022) proposed any-order modeling, which allows a model to generate in any desired order. Shih et al. (2022) made the same observation that any-order models by default have to solve exponentially more masking problems than autoregressive models. However, whereas our work shows that learning in the face of this challenging task diversity can benefit the model at inference time, their work sought to alleviate complexity at training time by reducing the number of masking problems that need to be solved.
@@ Appendix C Experimental details in Section 3
As explained in Section 3.2, to evaluate the scaling law of the $\pi$-learner, we can simply adapt the autoregressive training setup (a transformer with causal attention) by modifying the input to $\pi(x_{0})$ and using a learnable positional embedding layer instead of RoPE. We borrow the training configurations from (Nie et al., 2024), which are also consistent with the TinyLlama (Zhang et al., 2024) configurations. In particular, we use AdamW optimizer (Loshchilov & Hutter, 2019), setting $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a weight decay of $0.1$ and $L=2048$. A cosine learning rate schedule is applied, with a maximum learning rate of $4\times 10^{-4}$ and a minimum learning rate of $4\times 10^{-5}$. We also note that unless otherwise specified, we maintain the same training configuration throughout the paper.
@@ D.1.2 Experiment on text data
Note that this approach has also been employed for unconditional sampling (Wang et al., 2024; Zheng et al., 2024b).
@@ Appendix E Omitted proofs
We first re-state the Proposition 3.1 from (Zheng et al., 2024a). To clarify, (Zheng et al., 2024a) generally considers the case beyond the time-embedding denoising network $p_{\theta}$.
Proposition E.1 (Proposition 3.1 of (Zheng et al., 2024a)).

### v2

@@ 1 Introduction
While diffusion models (Ho et al., 2020; Song et al., 2021) are now the dominant approach for generative modeling in continuous domains like image, video, and audio, efforts to extend this methodology to discrete domains like text and proteins (Austin et al., 2021; Lou et al., 2024; Hoogeboom et al., 2021b) remain nascent. Among numerous proposals, masked diffusion models (MDMs) (Lou et al., 2024; Sahoo et al., 2025; Shi et al., 2024) have emerged as a leading variant, distinguished by a simple and principled objective: to generate samples, learn to reverse a noise process which independently and randomly masks tokens.
In many applications, such as language modeling, masked diffusion models (MDMs) still underperform compared to autoregressive models (ARMs) (Nie et al., 2024; Zheng et al., 2024), which instead learn to reverse a noise process that unmasks tokens sequentially from left to right. However, recent studies suggest that MDMs may offer advantages in areas where ARMs fall short, including reasoning (Nie et al., 2024; Kitouni et al., 2025), planning (Ye et al., 2024), and infilling (Gong et al., 2024). This raises a key question: what are the strengths and limitations of MDMs compared to ARMs, and under what conditions can MDMs be scaled to challenge the dominance of ARMs in discrete generative modeling?
@@
  • Flexibility at inference time: On the other hand, the sampling paths taken by an MDM are less rigid. The order in which tokens are decoded at inference time is random instead of fixed to left-to-right. In fact, even more is possible: MDMs can actually be used to decode in any order (Zheng et al., 2024).
@@
While the above might appear to be bad news for MDMs, in the second part of this paper we answer our guiding question in the affirmative by building upon the observation (Zheng et al., 2024) that MDMs which can perfectly solve all masking subproblems can be used to decode in any order.
@@ 2 Masked Diffusion Models (MDM)
In this section, we explain the framework of Masked Diffusion Models (Shi et al., 2024; Sahoo et al., 2025) and its interpretation as an order-agnostic learner. MDMs gradually add noise to the true discrete data and learn the marginal distribution of the induced reverse process. Below, we formulate the forward and reverse processes for MDMs.
@@
The reverse transition probability $q_{s|t}(x_{s}^{i}|x_{t},x_{0})$ is approximated using $g_{\theta}(x_{s}^{i}|x_{t})\triangleq q_{s|t}(x_{s}^{i}\,\lvert\,x_{t},x_{0}\leftarrow p_{\theta}(x_{t},t))$ where $p_{\theta}(x_{t},t)$ is a denoising network trained to predict the marginal on $x_{0}$ via an ELBO-based loss. To be precise, $q_{s|t}\left(x_{s}^{i}\mid x_{t},x_{0}\leftarrow p_{\theta}(x_{t},t)\right)$ indicates the conditional probability where $p_{\theta}(x_{t},t)$ is placed in the position of $x_{0}$ within $q_{s|t}(x_{s}^{i}\mid x_{t},x_{0})$.
@@ 2.1.1 Order-agnostic training of MDMs
Recent works (Zheng et al., 2024; Ou et al., 2024) have observed that the learning problem of MDM is equivalent to a masked language model. Building upon their analysis, we reformulate the loss $\mathcal{L}_{\theta}$ to show that $\mathcal{L}_{\theta}$ is a linear combination of the loss for all possible infilling masks. We first define $x_{0}[M]$ as a masked sequence, obtained from original sequence $x_{0}$ where indices in the mask set $M$ (regarded as a subset of $[L]\triangleq\{1,2,\ldots,L\}$) are replaced with mask token 00.
@@ 2.1.2 Order-agnostic inference of MDMs
The MDM inference can be decomposed into two steps: (a) randomly selecting a set of positions to unmask and (b) assigning token values to each position via the denoising network $p_{\theta}$. More precisely, we can reformulate the reverse process $x_{s}\sim g_{\theta}(\cdot|x_{t})$ as follows.
[…]
@@ 3.2 Empirical evidence of hardness via likelihoods
Recent studies (Nie et al., 2024; Zheng et al., 2024) have shown that masked diffusion models (MDMs) underperform compared to autoregressive models (ARMs) on natural text data. In this section, we provide evidence that this performance gap is primarily due to the order-agnostic training of MDMs. Since natural text follows a left-to-right token order, we demonstrate that as training deviates from this order, model performance gradually deteriorates.
@@ 4.1 Effective design of ordering oracle
Top-$K$ probability (Zheng et al., 2023). Suppose we want to unmask $K$ positions at time step $t$, i.e., select $|S|=K$. In the Top-$K$ strategy, the uncertainty of a position is estimated by the maximum probability assigned to any value in the vocabulary. More precisely, the certainty at position $i$ is $\max_{j\in\{0,\ldots,m-1\}}p_{\theta}(x^{i}=j|x_{t})$ and $\mathcal{F}(\theta,x_{t})=\text{Top }K\left(\max p_{\theta}(x^{i}|x_{t})\right)$.
Top-$K$ strategy is a good proxy for many tasks and works well in practice (Zheng et al., 2023; Ye et al., 2024; Wang et al., 2024). However, this approach can often provide misleading estimates of uncertainty. Consider when an MDM is confused between two token values, thus assigning them almost equal but high probabilities. In this case, Top-$K$ strategy may still choose to unmask this position, despite its uncertainty. To mitigate this issue, we propose the following alternative strategy.
@@ 4.4 Easy to hard generalization
We keep the training dataset the same as proposed in Shah et al. (2024). Shah et al. (2024) created this dataset from Radcliffe (2020) by selecting the puzzles that can be solved using 7 fixed strategies and do not require backtracking-based search. We use the remaining puzzles in Radcliffe (2020) as our hard dataset. Hence, these puzzles all use a strategy not seen during training and/or backtracking to obtain the correct solution.
@@
Method ##\##Param Accuracy
@@ Appendix A Related works
(Continuous) diffusion models were originally built on continuous-space Markov chains with Gaussian transition kernels (Sohl-Dickstein et al., 2015; Ho et al., 2020). This was later extended to continuous time through the theory of stochastic differential equations (Song et al., 2021). In a similar vein, discrete diffusion models have emerged from discrete-space Markov chains (Hoogeboom et al., 2021b). Specifically, (Austin et al., 2021) introduced D3PM with various types of transition matrices. Later, Lou et al. (2024) proposed SEDD, incorporating a theoretically and practically robust score-entropy objective. Additionally, Varma et al. (2024); Liu et al. (2024b) introduced novel modeling strategies that classify tokens in a noisy sequence as either signal (coming from clean data) or noise (arising from the forward process). In particular, Liu et al. (2024b) uses this to give a planner that adaptively determines which tokens to denoise. While this is similar in spirit to our general discussion about devising adaptive inference strategies, we emphasize that their approach is specific to discrete diffusions for which the forward process scrambles the token values, rather than masking them.
@@
Meanwhile, the absorbing transition kernel has gained popularity as a common choice due to its better performance than other kernels. Building on this, Sahoo et al. (2025); Shi et al. (2024) aligned its framework with continuous diffusion, resulting in a simple and principled training recipe, referring to it as Masked Diffusion Model. Subsequent studies have explored various aspects of MDM. Gong et al. (2024) efficiently trained MDM via adaptation from autoregressive models, scaling MDM up to 7B parameters. Zheng et al. (2024) interpreted MDMs as order-agnostic learners and proposed a first-hitting sampler based on this insight. Ye et al. (2024); Gong et al. (2024) demonstrated that MDM outperforms autoregressive models in reasoning and planning tasks, emphasizing its impact on downstream applications. Nie et al. (2024) examined the scaling laws of MDM, while Xu et al. (2024); Liu et al. (2024a) identified limitations in capturing coordinate dependencies when the number of sampling steps is small and proposed additional modeling strategies to address this issue. Schiff et al. (2024) studied conditional generation using MDM and Rector-Brooks et al. (2024) tackled the challenge of controlling generated data distributions through steering methodologies. Chen & Ying (2024) provided a theoretical analysis showing that sampling error is small given accurate score function estimation.
@@
Even though language tasks generally have a natural order of “left-to-right” token generation, in many tasks like planning, reasoning, and combinatorial optimization, the natural order of token generation can be quite different from “left-to-right”. Even though prominent autoregressive-based language models achieve impressive performance on various tasks, many works (Golovneva et al., 2024; Chen et al., 2024; Kitouni et al., 2025) have shown that this performance is tied to the training order of the tasks and therefore can cause brittleness from it. For example, Chen et al. (2024) showed that simply permuting the premise order on math tasks causes a performance drop of 30%. The reason behind such brittleness regarding the ordering is the inherent “left-to-right” nature of the autoregressive models. Several works (Liao et al., 2020) have tried to address this issue in the autoregressive framework. In particular, (Papadopoulos et al., 2024) highlighted the significance of left-to-right ordering in natural language by comparing its likelihood to that of the reverse (right-to-left) ordering.
Recently, discrete diffusion models have emerged as a promising approach for discrete data apart from autoregressive models. Additionally, the order-agnostic training of discrete diffusion models opens up the multiple sampling paths during the inference but it also faces some challenges during the training therefore, they seem a promising approach to elicit any order reasoning. Zheng et al. (2023) proposed different ways of implementing an adaptive inference strategy for MDM but a concrete understanding of why such an adaptive inference strategy is needed is still lacking. In this work, we explore various aspects of vanilla MDM training and how adaptive MDM inference can mitigate the issues raised by vanilla MDM training and elicit any order reasoning.
@@
Efforts to learn the natural language using non-autoregressive modeling began with BERT (Devlin et al., 2019). Non-causal approaches can take advantage of the understanding the text data representation. (Chang et al., 2022) adopted a similar approach for learning image representations. Building on these intuitions, (Shih et al., 2022; Hoogeboom et al., 2021a) proposed any-order modeling, which allows a model to generate in any desired order. Shih et al. (2022) made the same observation that any-order models by default have to solve exponentially more masking problems than autoregressive models. However, whereas our work shows that learning in the face of this challenging task diversity can benefit the model at inference time, their work sought to alleviate complexity at training time by reducing the number of masking problems that need to be solved.
@@ Appendix C Experimental details in Section 3
As explained in Section 3.2, to evaluate the scaling law of the $\pi$-learner, we can simply adapt the autoregressive training setup (a transformer with causal attention) by modifying the input to $\pi(x_{0})$ and using a learnable positional embedding layer instead of RoPE. We borrow the training configurations from (Nie et al., 2024), which are also consistent with the TinyLlama (Zhang et al., 2024) configurations. In particular, we use AdamW optimizer (Loshchilov & Hutter, 2017), setting $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a weight decay of $0.1$ and $L=2048$. A cosine learning rate schedule is applied, with a maximum learning rate of $4\times 10^{-4}$ and a minimum learning rate of $4\times 10^{-5}$. We also note that unless otherwise specified, we maintain the same training configuration throughout the paper.
@@ D.1.2 Experiment on text data
Note that this approach has also been employed for unconditional sampling (Wang et al., 2024; Zheng et al., 2023).
@@ Appendix E Omitted proofs
We first re-state the Proposition 3.1 from (Zheng et al., 2024). To clarify, (Zheng et al., 2024) generally considers the case beyond the time-embedding denoising network $p_{\theta}$.
Proposition E.1 (Proposition 3.1 of (Zheng et al., 2024)).